Let (V, <, >) be an inner product space and

let T belongs to A(V). Prove that the following 

conditions are equivalent. 

(i) T*T =I

(ii) <Tx, Ty> = <x,y> for all x, yEV 

(iii) ||Tx||=||x||for all xEV. 


1
Expert's answer
2021-01-05T13:56:26-0500

"(iii)\rightarrow (ii)"

Suppose that "||T(x)|| = ||x||" for all "x \u2208 V".

Now for any "x, y ∈ V :\;\; <x,y>=\frac{1}{4}\sum_{k=1}^{4}i^k||x+i^ky||^2"

if "\mathbb{F}=\mathbb{C},"

And if "\mathbb{F}=\R" ,

we have similar result: "<x,y>=\frac{1}{4}\sum_{k=1}^{4}(||x+y||^2-||x-y||^2)"

Using the above relation for the complex version, we have

"<T(x),T(y)>\\=\frac{1}{4}\sum_{k=1}^{4}i^k||T(x)+i^kT(y)||^2\\
=\frac{1}{4}\sum_{k=1}^{4}i^k||T(x+i^ky)||^2\\=\frac{1}{4}\sum_{k=1}^{4}i^k||x+i^ky||^2\\
=<x,y>"

"(ii)\rightarrow (iii)"

Suppose that "<T(x),T(y)>=<x,y>\forall x,y\in V".

Substituting "y = x," we have the desired result. 


"(ii)\rightarrow (i)"

Suppose "T" is unitary.

Then "T" has an inverse "T^{-1}".

So, for "x, y \u2208 V" ,

"<T(x), y> = <T(x), TT^ {\u22121} (y)> = <x, T^{\u22121} (y)>" .

So, "T^*" exists and "T^*=T^{-1}"

"\Rightarrow TT^*=I"


"(i)\rightarrow (ii)"

Assume the adjoint "T^*" exists

and "TT^*=T^*T=I"

Now for "x, y \u2208 V", we have "<T(x), T(y)) = <x, T^ \u2217T(y)> = <x, y>"


Therefore from above discussion, we can conclude that,

"(iii)\iff(ii)\;\; and \;\;(i)\iff (ii)"

So that implies "(i)\iff(ii)\iff(iii)"

"\therefore" "\pmb{(i),(ii),(iii)}" are equivalent.